a line has a slope of -2/3 and y-intercept of (0,3) what is the value of y when x=4
Answer by graphmatics(170)

We use the general equation y-y1=m(x-x1) where m is the slope and (x1,y1) is apoint on the line. So
y-3=(-2/3)(x-0)
y=(-2/3)x+3
If x=4 then y=-8/3+3=-8/3+9/3=1/3